OpenStudy (anonymous):

1.The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (agent0smith):

You want to get a by itself... d = ut + 0.5at2 start by subtracting ut from both sides THEN multiply both sides by 2 THEN divide both sides by t^2

OpenStudy (agent0smith):

d = ut + 0.5at^2 subtract ut from both sides to get d - ut = 0.5at^2

OpenStudy (agent0smith):

Try the next two steps on your own.

OpenStudy (anonymous):

I do not understand that.

OpenStudy (anonymous):

@madrockz

OpenStudy (anonymous):

d=ut+0.5at^2 u have to find the formula for a.so u have to keep a alone on one side and free from all the coefficients. 0.5at^2+ut=d 0.5at^2=d-ut 0.5a=(d-ut)/t^2

OpenStudy (anonymous):

now can u find the value of a?

OpenStudy (anonymous):

@BIGDOG96

OpenStudy (anonymous):

@redred413 @CGGURUMANJUNATH are you going to help, or just look at it ?

OpenStudy (cggurumanjunath):